OpenStudy (anonymous):

find the missing dimension of a circle. the area is 81(pi)ft square

OpenStudy (anonymous):

(using A=3.14r^2) 81 multiplied by 3.14=254.34 254.34=3.14r^2 divide both sides by 3.14 to isolate the "r" 81=r^2 then ... 9=r